find the radius of the circle :
a)when theta=90 ,A=20cm^2 ?
NOTE:A is chord area ?

please helpppp ?

Answer by rothauserc(4718)

The chord area = (R^2/2) * ( ((theta * pi)/180) - sin theta), where R is the radius
:
Note the above says the chord area = area of circular sector - area of triangular portion
:
we are given the chord area = 20 cm^2 and theta = 90
:
20 = (R^2/2) * ( ((90 * pi)/180) - sin 90)
:
(R^2/2) * ( (pi/2) - 1) = 20
:
(R^2/2) * 0.5708 = 20
:
(R^2/2) = 35.0385
:
R^2 = 70.077
:
*************************************************************************
Area of circle = pi * R^2 = pi * 70.077 = 220.1534 approximately 220 cm^2
*************************************************************************
: